OpenStudy (anonymous):

Solve. 36=5/9k PLEASE HELP!!!

OpenStudy (pinksapphire):

What is the question? Are you solving for k?

OpenStudy (anonymous):

yes the choices are A. 4/5 B. 20 C. 60 and 4/5 D. 180

OpenStudy (pinksapphire):

First, multiply by the denominator: \[9(36)=(\frac{ 5 }{ 9 }k)9\]

OpenStudy (anonymous):

ok whats the next step?

OpenStudy (pinksapphire):

\[\frac{ 324 }{ 5 }=\frac{ 5k }{ 5 }\]

OpenStudy (pinksapphire):

The answer will be 64.8

OpenStudy (anonymous):

thanks so C?

OpenStudy (pinksapphire):

C is 60.8 right?

OpenStudy (pinksapphire):

Because the answer that I got was 64.8 and if you plug it in for "k" you will get 36=36...